Consider the following observations on a receptor binding measure (adjusted distribution volume) for a sample of 13 healthy individuals: 23, 39, 40, 41, 44, 48, 52, 58, 64, 67, 68, 69, 71. (a) Is it plausible that the population distribution from which this sample was selected is normal? [ Yes , it ( is +) plausible that the population distribution is normal. (b) Calculate an interval for which you can be 95% confident that at least 95% of all healthy individuals in the population have adjusted distribution volumes ying between the limits of the interval. (Round your answers to three decimal places.) 44.520x61.217X c) Predict the adjusted distribution volume of a single healthy individual by calculating a 95% prediction interval Round your answers to three decimal places. How does this intervals width compare to the width of the interval calculated in part (b)? This intervals width is les than the width of the interval calculated in part b). You may need to use the appropriate table in the Appendix of Tables to answer this question Need Help? dtTalk to a Tuterpart b and c!!! please help!
Answer #1

b)

sample mean         x= 52.615
sample size             n= 13.000
sample std deviation s= 14.8915
std error sx=s/√n= 4.1302
for 95% CI; and 12 df, critical t= 2.1790
margin of error E=t*std error                            = 9.000
lower bound=sample mean-E = 43.616
Upper bound=sample mean+E= 61.615
from above 95% confidence interval for population mean =(43.616 ,61.615)

c)

std errror of mean ='s=s*√(1+1/n)= 15.454
for 95% CI; and 12 df, value of t= 2.1790
margin of error E=t*std error     =                                                    33.6735
lower bound=sample mean-margin of error = 18.942
Upper bound=sample mean +margin of error= 86.289

95% prediction interval for population mean =(18.942, 86.289)